# The role of turbulence in setting the phase of the ISM and implications for the star formation rate

### v1

@@ Acknowledgements
This research has received funding from the European Research Council synergy grant ECOGAL (Grant 855130). TC would like to thank all the members of the ECOGAL collaboration for discussions and comments which greatly improved this work. This work also benefited from the Grand Cascade workshop hosted at Institute Pascal in Orsay, France. This work was granted access to HPC resources of CINES and CCRT under the allocation x2020047023 made by GENCI (Grand Equipement National de Calcul Intensif). RSK and SCOG acknowledge funding from the German Research Foundation (DFG) under Germany’s Excellence Strategy for the STRUCTURES cluster of excellence in Heidelberg (EXC-2181/1 - 390900948) and also via the collaborative research center (SFB 881, Project-ID 138713538) “The Milky Way System” (subprojects A1, B1, B2, and B8).

### v2

@@ Acknowledgements
This research has received funding from the European Research Council synergy grant ECOGAL (Grant 855130). TC would like to thank all the members of the ECOGAL collaboration for discussions and comments which greatly improved this work. The authors acknowledge Interstellar Institute’s programs ”II4: The Grand Cascade”, as well as the Paris-Saclay University’s Institut Pascal for hosting discussions that triggered this work and nourished the development of the ideas behind it. This work was granted access to HPC resources of CINES and CCRT under the allocation x2020047023 made by GENCI (Grand Equipement National de Calcul Intensif). NB acknowledges support from the ANR BRIDGES grant (ANR-23-CE31-0005). In addition to ECOGAL, RSK and SCOG acknowledge funding from the German Excellence Strategy via the Heidelberg Excellence Cluster STRUCTURES (EXC-2181/1 - 390900948). TC acknowledges funding received from the European High Per- formance Computing Joint Undertaking (JU) and Belgium, Czech Republic, France, Germany, Greece, Italy, Norway, and Spain under grant agreement no. 101093441 (SPACE).
@@ Appendix A WNM temperatures for different ionisation fractions
Appendix B The influence of the supernova treatment
Figure 19: Comparisons between simulations for which the supernova velocities are limited to 300 km s-1 (dashed lines) and simulations where they are limited to 3000 km s-1 (solid lines). The left panel portrays the volume weighted density PDF and the right panel the mean velocity in density bins.
As explained above, the time step induced by supernova explosions is rather short. For instance, with a resolution of 4 pc it is about 100 yr or less, making the simulations very timely, even with a relatively coarse spatial resolution. For a sub-parsec resolution, as we have in this work, this would produce time steps on the order of 10 yr or even less. To circumvent this difficulty, we have limited the velocity of the supernova remnant to 300 km s-1 and the temperature of the gas to 106 K. With these limits, the time steps is about 10 times larger. However, it is important to check the effects that this procedure may have on the results. Figure 19 compares the results of simulations where the velocities and temperatures use the limitations listed in Section 2 with those of simulations for which the limits are set to much higher values (3000 km s-1 and 108 K). On the left, we show the density PDF and on the right is the velocity dispersion within density bins. The column density of these two simulations is 19 M pc-2. As can be seen, large differences are found for the diffuse gas (i.e. below  0.1 cm-3), for both the density PDF and the velocity dispersion. However, the differences for gas denser than  0.1 cm-3 appear to be quite limited. Let us recall that for the spatial scales at which ISM simulations are run, large velocities appear only when supernovae explode in very diffuse gas (i.e. below 0.1 cm-3). In this case, the supernova explosion has little impact on the dense gas.